# Photo-SLAM: Real-time Simultaneous Localization and Photorealistic Mapping for Monocular, Stereo, and RGB-D Cameras

### 28 Nov 2023 v1

@@ Abstract
The integration of neural rendering and the SLAM system recently showed promising results in joint localization and photorealistic view reconstruction. However, existing methods, fully relying on implicit representations, are so resource-hungry that they cannot run on portable devices, which deviates from the original intention of SLAM. In this paper, we present Photo-SLAM, a novel SLAM framework with a hyper primitives map. Specifically, we simultaneously exploit explicit geometric features for localization and learn implicit photometric features to represent the texture information of the observed environment. In addition to actively densifying hyper primitives based on geometric features, we further introduce a Gaussian-Pyramid-based training method to progressively learn multi-level features, enhancing photorealistic mapping performance. The extensive experiments with monocular, stereo, and RGB-D datasets prove that our proposed system Photo-SLAM significantly outperforms current state-of-the-art SLAM systems for online photorealistic mapping, e.g., PSNR is 30% higher and rendering speed is hundreds of times faster in the Replica dataset. Moreover, the Photo-SLAM can run at real-time speed using an embedded platform such as Jetson AGX Orin, showing the potential of robotics applications.
@@ 1 Introduction
Figure 1: Photo-SLAM is a novel framework for simultaneous localization and photorealistic mapping in real-time supporting monocular, stereo, and RGB-D cameras. It can reconstruct high-fidelity views of scenes while render speed is up to 1000 FPS.
@@
Despite the promising results achieved through the integration of neural rendering and SLAM, existing methods simply and heavily rely on implicit representations, making them computationally intensive and unsuitable for deployment on resource-constrained devices. For example, Nice-SLAM [45] leverages a hierarchical grid [41] to store learnable features representing the environment while ESLAM [16] utilizes multi-scale compact tensor components [3]. They then jointly estimate the camera poses and optimize features by minimizing the reconstruction loss of a batch of ray sampling [21]. Such an optimization process is time-consuming. Consequently, it is indispensable for them to incorporate corresponding depth information obtained from various sources such as RGB-D cameras, dense optical flow estimators [32], or monocular depth estimators [12] to ensure efficient convergence. Additionally, since the implicit features are decoded by the multi-layer perceptrons (MLP), it is typically necessary to carefully define a bounding area to normalize ray sampling for optimal performance [14]. It essentially limits the scalability of the system. These limitations imply that they cannot provide real-time exploration and mapping capabilities in the unknown environment using portable platforms, which is one of the main objectives of SLAM.
@@
  • We developed the first simultaneous localization and photorealistic mapping system based on hyper primitives map. The novel framework supports monocular, stereo, and RGB-D cameras, in indoor and outdoor environments.
@@
  • The system, fully implemented in C++ and CUDA, achieves start-of-the-art performance and can run at real-time speed even on embedded platforms. The code will be publicly available.
@@ 3.3 Photorealisitc Mapping
$\mathcal{L}=(1-\lambda)\left|{I_{\text{r}}-I_{\text{gt}}}\right|_{1}+\lambda\frac{1-\text{SSIM}(I_{\text{r}},I_{\text{gt}})}{2},$ (4)
@@ 4 Experiment
Cam Method RMSE $\downarrow$ STD $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Tracking FPS $\uparrow$ Rendering FPS $\uparrow$ GPU Memory Usage $\downarrow$
Mono ORB-SLAM3 [2] 3.942 3.115 - - - 58.749 - 0
DROID-SLAM [33] 0.725 0.308 - - - 35.473 - 11 GB
Nice-SLAM* [45] 99.9415 35.336 16.311 0.720 0.439 2.384 0.944 12 GB
Orbeez-SLAM [4] - - 23.246 0.790 0.336 49.200 1.030 6 GB
Go-SLAM [43] 71.054 24.593 21.172 0.703 0.421 25.366 0.821 22 GB
Ours (Jetson) 1.235 0.756 29.284 0.883 0.139 18.315 95.057 4 GB
Ours (Laptop) 0.713 0.524 33.049 0.926 0.086 19.974 353.504 4 GB
Ours 1.091 0.892 33.302 0.926 0.078 41.646 911.262 6 GB
RGB-D ORB-SLAM3 [2] 1.833 1.478 - - - 52.209 - 0
DROID-SLAM [33] 0.634 0.248 - - - 36.452 - 11 GB
BundleFusion [6] 1.606 0.969 23.839 0.822 0.197 8.630 - 5 GB
Nice-SLAM [45] 2.350 1.590 26.158 0.832 0.232 2.331 0.611 12 GB
Orbeez-SLAM [4] 0.888 0.562 32.516 0.916 0.112 41.333 1.401 6 GB
ESLAM [16] 0.568 0.274 30.594 0.866 0.162 6.687 2.626 21 GB
Co-SLAM [35] 1.158 0.602 30.246 0.864 0.175 14.575 3.745 4 GB
Go-SLAM [43] 0.571 0.218 24.158 0.766 0.352 19.437 0.444 24 GB
Ours (Jetson) 0.581 0.289 31.978 0.916 0.101 17.926 116.395 4 GB
Ours (Laptop) 0.590 0.289 34.853 0.944 0.062 20.597 396.082 4 GB
Ours 0.604 0.298 34.958 0.942 0.059 42.485 1084.017 5 GB
@@
Figure 7: Mapping comparison with other monocular camera systems.
@@
4.1 Implementation and Experiment Setup
We implemented Photo-SLAM fully in C++ and CUDA, making use of ORB-SLAM3 [2], 3D Gaussian splatting [18], and the LibTorch framework. The optimization of photorealistic mapping is performed with the Stochastic Gradient Descent algorithm while we use a fixed learning rate and $\lambda=0.2$. Considering image resolution of the testing dataset, the level of the Gaussian pyramid is set to three, i.e. $n=2$ by default. The compared baseline includes a SOTA classical SLAM system ORB-SLAM3 [2], a real-time RGB-D dense reconstruction system BundleFusion [6], a deep-learning based system DROID-SLAM [33], and recent SLAM systems supporting view synthesis, i.e. Nice-SLAM [45], Orbeez-SLAM [4], ESLAM [16], Co-SLAM [35], and Go-SLAM [43].

### 08 Apr 2024 v2

@@ Abstract
The integration of neural rendering and the SLAM system recently showed promising results in joint localization and photorealistic view reconstruction. However, existing methods, fully relying on implicit representations, are so resource-hungry that they cannot run on portable devices, which deviates from the original intention of SLAM. In this paper, we present Photo-SLAM, a novel SLAM framework with a hyper primitives map. Specifically, we simultaneously exploit explicit geometric features for localization and learn implicit photometric features to represent the texture information of the observed environment. In addition to actively densifying hyper primitives based on geometric features, we further introduce a Gaussian-Pyramid-based training method to progressively learn multi-level features, enhancing photorealistic mapping performance. The extensive experiments with monocular, stereo, and RGB-D datasets prove that our proposed system Photo-SLAM significantly outperforms current state-of-the-art SLAM systems for online photorealistic mapping, e.g., PSNR is 30% higher and rendering speed is hundreds of times faster in the Replica dataset. Moreover, the Photo-SLAM can run at real-time speed using an embedded platform such as Jetson AGX Orin, showing the potential of robotics applications. Project Page and code: https://huajianup.github.io/research/Photo-SLAM/.
@@ 1 Introduction
Figure 1: Rendering and trajectory results. Photo-SLAM can reconstruct high-fidelity views of scenes using monocular, stereo, and RGB-D cameras while render speed is up to 1000 FPS.
@@
Despite the promising results achieved through the integration of neural rendering and SLAM, existing methods simply and heavily rely on implicit representations, making them computationally intensive and unsuitable for deployment on resource-constrained devices. For example, Nice-SLAM [46] leverages a hierarchical grid [42] to store learnable features representing the environment while ESLAM [16] utilizes multi-scale compact tensor components [3]. They then jointly estimate the camera poses and optimize features by minimizing the reconstruction loss of a batch of ray sampling [21]. Such an optimization process is time-consuming. Consequently, it is indispensable for them to incorporate corresponding depth information obtained from various sources such as RGB-D cameras, dense optical flow estimators [33], or monocular depth estimators [12] to ensure efficient convergence. Additionally, since the implicit features are decoded by the multi-layer perceptrons (MLPs), it is typically necessary to carefully define a bounding area to normalize ray sampling for optimal performance, as discussed in [14]. It essentially limits the scalability of the system. These limitations imply that they cannot provide real-time exploration and mapping capabilities in the unknown environment using portable platforms, which is one of the main objectives of SLAM.
@@
  • We developed the first simultaneous localization and photorealistic mapping system based on hyper primitives map. The novel framework supports monocular, stereo, and RGB-D cameras in indoor and outdoor environments.
@@
  • The system, fully implemented in C++ and CUDA, achieves start-of-the-art performance and can run at real-time speed even on embedded platforms.
@@ 3.3 Photorealisitc Mapping
$\mathcal{L}=(1-\lambda)\left|{I_{\text{r}}-I_{\text{gt}}}\right|_{1}+\lambda{(1-\text{SSIM}(I_{\text{r}},I_{\text{gt}}))},$ (4)
@@ 4 Experiment
Cam Method RMSE $\downarrow$ STD $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Operation Time $\downarrow$ Tracking FPS $\uparrow$ Rendering FPS $\uparrow$ GPU Memory Usage $\downarrow$
Mono ORB-SLAM3 [2] 3.942 3.115 - - - $<$1 mins 58.749 - 0
DROID-SLAM [34] 0.725 0.308 - - - $<$2 mins 35.473 - 11 GB
Nice-SLAM* [46] 99.9415 35.336 16.311 0.720 0.439 $>$10 mins 2.384 0.944 12 GB
Orbeez-SLAM [4] - - 23.246 0.790 0.336 $<$5 mins 49.200 1.030 6 GB
Go-SLAM [44] 71.054 24.593 21.172 0.703 0.421 $<$5 mins 25.366 0.821 22 GB
Ours (Jetson) 1.235 0.756 29.284 0.883 0.139 $<$5 mins 18.315 95.057 4 GB
Ours (Laptop) 0.713 0.524 33.049 0.926 0.086 $<$5 mins 19.974 353.504 4 GB
Ours 1.091 0.892 33.302 0.926 0.078 $<$2 mins 41.646 911.262 6 GB
RGB-D ORB-SLAM3 [2] 1.833 1.478 - - - $<$1 mins 52.209 - 0
DROID-SLAM [34] 0.634 0.248 - - - $<$2 mins 36.452 - 11 GB
BundleFusion [6] 1.606 0.969 23.839 0.822 0.197 $<$5 mins 8.630 - 5 GB
Nice-SLAM [46] 2.350 1.590 26.158 0.832 0.232 $>$10 mins 2.331 0.611 12 GB
Orbeez-SLAM [4] 0.888 0.562 32.516 0.916 0.112 $<$5 mins 41.333 1.401 6 GB
ESLAM [16] 0.568 0.274 30.594 0.866 0.162 $<$5 mins 6.687 2.626 21 GB
Co-SLAM [36] 1.158 0.602 30.246 0.864 0.175 $<$5 mins 14.575 3.745 4 GB
Go-SLAM [44] 0.571 0.218 24.158 0.766 0.352 $<$5 mins 19.437 0.444 24 GB
Point-SLAM [27] 0.596 0.249 34.632 0.927 0.083 $>$2 hrs 0.345 0.510 24 GB
Ours (Jetson) 0.581 0.289 31.978 0.916 0.101 $<$5 mins 17.926 116.395 4 GB
Ours (Laptop) 0.590 0.289 34.853 0.944 0.062 $<$5 mins 20.597 396.082 4 GB
Ours 0.604 0.298 34.958 0.942 0.059 $<$2 mins 42.485 1084.017 5 GB
@@
Figure 7: Mapping comparison with other monocular camera systems on Replica office3 scene.
@@
4.1 Implementation and Experiment Setup
We implemented Photo-SLAM fully in C++ and CUDA, making use of ORB-SLAM3 [2], 3D Gaussian splatting [18], and the LibTorch framework. The optimization of photorealistic mapping is performed with the Stochastic Gradient Descent algorithm while we use a fixed learning rate and $\lambda=0.2$. Considering image resolution of the testing dataset, the level of the Gaussian pyramid is set to three, i.e. $n=2$ by default. The compared baseline includes a SOTA classical SLAM system ORB-SLAM3 [2], a real-time RGB-D dense reconstruction system BundleFusion [6], a deep-learning based system DROID-SLAM [34], and recent SLAM systems supporting view synthesis, i.e. Nice-SLAM [46], Orbeez-SLAM [4], ESLAM [16], Co-SLAM [36], and Point-SLAM [27] and Go-SLAM [44].
4.2 Results and Evaluation
@@ 5 Conclusion
Acknowledgements: This work is partially supported by the Innovation and Technology Support Programme of the Innovation and Technology Fund (Ref: ITS/200/20FP), the Marine Conservation Enhancement Fund (MCEF20107 & MCEF23EG01), and an internal grant from HKUST (R9429).
Supplementary Material
Photo-SLAM is a novel system for simultaneous localization and photorealistic mapping, which can even run on embedded platforms at real-time speed, as demonstrated in Fig. 11. In this supplementary, we provide additional results regarding localization and mapping performance.
Figure 11: The proposed system has real-time performance on embedded platforms, such as Jetson AGX Orin Developer Kit.
Figure 12: Tracking speed comparisons on scene office0 using an RGB-D camera. The vertical axis denotes the processing time of each frame while the horizontal axis denotes the frame number. [Ave/Min/Max/Std] represent the average, minimum, and maximum tracking time and its standard deviation respectively.
6 Localication
Stability. As online systems, SLAMs are required to process the incoming frames and estimate current camera poses in time. Therefore, tracking stability regarding latency and the average processing time is an important factor in evaluating system performance in addition to pose estimation accuracy. As reported in Table 1 of the main paper, Photo-SLAM is capable of processing more than 40 frames per second with accurate pose estimation. The average tracking speed is about six times faster than ESLAM [16] and three times faster than Co-SLAM [36]. Here, we provide additional analysis on tracking stability while an example plotted in Fig. 12.
Although the average tracking time of Go-SLAM [44] is less than Co-SLAM and ESLAM, the processing latency is high due to frequently conducting expensive global optimization. As shown in Fig. 12, Go-SLAM often takes about 1 second to process the frame and estimate the pose. Moreover, both Nice-SLAM [46] and Co-SLAM need a longer time to accurately initialize the tracking. Obviously, our system can rapidly and stably process the incoming frames, having minimum average tracking time and standard deviation. The peak processing time of our system occurs when loop closure is detected for correcting pose estimation drift.
Accuracy. Some qualitative tracking results of Photo-SLAM are demonstrated in Fig. 13.
Figure 13: Trajectory in the reconstructed map. Green points denote ground truth trajectory while red denotes the estimated trajectory of Photo-SLAM.
7 Discussion
Online Mapping vs Offline Mapping. For online mapping, the mapping process occurs simultaneously with the localization process. Therefore it requires continuous and prompt updates with each new observation as the robot or camera moves and observes its surroundings. In general, online photorealistic mapping is more challenging than offline photorealistic mapping, since it is crucial to balance the trade-off between computational efficiency and rendering quality. As mentioned in the main paper, we proposed a geometry-based densification strategy and a Gaussian-Pyramid-based (GP) learning method to achieve high-quality online mapping. To further support this statement, we compared the photorealistic mapping performance between our Photo-SLAM and 3D Gaussian splatting (3DG) [18]. 3DG is the SOTA offline method which takes a set of images with known poses and a sparse point cloud as input to learn a radiance field for view synthesis. During the experiments, 3DG used the keyframe poses estimated by Photo-SLAM and performed training for the same duration as Photo-SLAM. The required point cloud input is initialized in three different ways: 1) randomly initializing 100 points; 2) randomly initializing 10,000 points; and 3) initializing from the hyper primitives map of Photo-SLAM. The results are reported in Table 5. Without inputting fine-grained point clouds, 3DG needs more time for optimization such that the rendering quality decreases. In addition, to enhance rendering quality, 3DG tends to densify point clouds leading to larger model size and slower rendering speed. Whether using monocular cameras or RGB-D cameras, Photo-SLAM consistently delivers compelling rendering quality and faster rendering speeds, owing to the effectiveness of the proposed algorithms.
method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Rendering FPS $\uparrow$ Model Size (MB)
1) 3DG 27.844 0.861 0.213 745.480 36.141
2) 3DG 34.555 0.942 0.065 483.904 144.196
3) 3DG 37.055 0.962 0.032 448.109 219.470
Ours using Mono 33.302 0.926 0.078 911.262 31.419
Ours using RGB-D 34.958 0.942 0.059 1084.017 35.211
Table 5: Comparison of mapping performance between 3D Gaussian splatting (3DG) and our system Photo-SLAM with different settings on the Replica dataset.
On TUM Dataset Resources
Scene Cam Method Tracking FPS $\uparrow$ Rendering FPS $\uparrow$ Model Size (MB)
fr1-desk Mono Ours (Jetson) 28.267 340.507 4.610
Ours (Laptop) 28.330 1105.062 7.421
Ours 57.781 2016.690 10.027
RGB-D Ours (Jetson) 27.970 380.622 5.743
Ours (Laptop) 28.930 1061.040 8.432
Ours 58.378 2083.896 9.963
fr2-xyz Mono Ours (Jetson) 24.005 169.321 14.286
Ours (Laptop) 24.922 619.554 16.102
Ours 58.241 1405.797 20.380
RGB-D Ours (Jetson) 21.032 274.718 6.319
Ours (Laptop) 22.665 701.590 13.850
Ours 52.904 1790.120 21.399
fr3-office Mono Ours (Jetson) 36.700 291.398 10.669
Ours (Laptop) 38.929 824.658 16.249
Ours 81.575 1522.120 19.211
RGB-D Ours (Jetson) 18.039 291.907 12.726
Ours (Laptop) 19.636 764.342 15.349
Ours 43.650 1540.757 17.009
Table 6: Additional results of Photo-SLAM with different platforms on the TUM dataset.
On EuRoC Dataset Resources
Scene Method Tracking FPS $\uparrow$ Rendering FPS $\uparrow$ Model Size (MB)
MH-01 Ours (Jetson) 21.359 93.762 43.385
Ours (Laptop) 25.019 316.403 89.700
Ours 44.977 613.958 123.528
MH-02 Ours (Jetson) 22.355 101.021 36.263
Ours (Laptop) 26.189 332.174 81.569
Ours 46.556 675.508 113.116
V1-01 Ours (Jetson) 21.332 106.008 28.444
Ours (Laptop) 25.403 367.903 55.263
Ours 44.763 835.119 74.457
V2-01 Ours (Jetson) 23.872 99.988 27.840
Ours (Laptop) 27.556 307.025 62.588
Ours 48.911 595.234 82.600
Table 7: Additional results of Photo-SLAM with different platforms on the EuRoC MAV stereo dataset.
On Replica Dataset Localization Mapping Resources
[…]
office0 Mono Ours (Jetson) 0.467 0.00334 34.415 0.940 0.949 18.328 123.551 17.703
Ours (Laptop) 0.587 0.00343 36.227 0.954 0.071 20.422 413.645 21.703
Ours 0.575 0.00369 36.989 0.955 0.061 42.487 930.598 24.975
RGB-D Ours (Jetson) 0.499 0.00356 35.447 0.949 0.086 19.076 154.087 18.281
Ours (Laptop) 0.519 0.00317 38.219 0.965 0.053 22.446 497.917 18.826
Ours 0.522 0.00307 38.477 0.964 0.050 48.588 1447.887 19.740
office1 Mono Ours (Jetson) 5.586 0.31140 32.382 0.904 0.113 18.312 80.048 22.904
Ours (Laptop) 0.379 0.00463 37.970 0.954 0.060 19.968 322.160 21.224
Ours 0.315 0.00383 37.592 0.950 0.062 42.296 857.324 26.982
RGB-D Ours (Jetson) 0.402 0.00517 37.510 0.953 0.065 19.194 118.584 20.656
Ours (Laptop) 0.440 0.00543 39.109 0.962 0.049 22.349 496.644 19.548
Ours 0.436 0.00477 39.089 0.961 0.047 47.333 1263.343 21.193
office2 Mono Ours (Jetson) 1.402 0.01452 28.083 0.900 0.131 17.502 90.181 19.704
Ours (Laptop) 2.087 0.02154 31.202 0.927 0.098 18.975 343.662 27.332
Ours 5.031 0.04696 31.794 0.929 0.091 39.604 930.777 31.558
RGB-D Ours (Jetson) 1.209 0.00964 29.755 0.919 0.110 17.860 124.420 27.560
Ours (Laptop) 1.188 0.00972 32.720 0.940 0.080 21.507 425.452 31.711
Ours 1.276 0.01094 33.034 0.938 0.077 44.062 904.249 34.065
office3 Mono Ours (Jetson) 0.429 0.00232 28.058 0.886 0.132 17.881 96.872 15.505
Ours (Laptop) 0.409 0.00239 32.012 0.924 0.090 19.518 368.530 20.475
Ours 0.472 0.00227 31.622 0.920 0.086 40.870 1131.957 26.653
RGB-D Ours (Jetson) 0.718 0.00222 30.954 0.917 0.103 17.889 120.118 20.270
Ours (Laptop) 0.747 0.00233 33.594 0.939 0.072 20.051 388.624 23.617
Ours 0.782 0.00233 33.789 0.938 0.066 40.603 1125.175 25.226
office4 Mono Ours (Jetson) 0.579 0.00305 30.399 0.921 0.109 18.755 102.949 15.201
Ours (Laptop) 0.616 0.00279 33.656 0.940 0.078 20.311 375.033 21.444
Ours 0.583 0.00272 34.168 0.941 0.072 42.262 849.305 26.154
RGB-D Ours (Jetson) 0.661 0.00367 32.219 0.931 0.091 17.107 92.237 32.405
Ours (Laptop) 0.629 0.00446 35.534 0.951 0.059 19.361 333.874 32.270
Ours 0.582 0.00423 36.020 0.952 0.054 39.870 1061.749 35.421
room0 Mono Ours (Jetson) 0.369 0.00321 26.423 0.787 0.221 17.987 87.246 17.121
Ours (Laptop) 0.349 0.00294 29.899 0.868 0.125 19.521 332.127 33.151
Ours 0.345 0.00299 29.772 0.871 0.106 41.020 754.729 44.333
RGB-D Ours (Jetson) 0.514 0.00265 27.867 0.833 0.165 17.424 104.248 31.196
Ours (Laptop) 0.521 0.00257 31.288 0.914 0.075 19.119 322.585 52.266
Ours 0.541 0.00270 30.716 0.899 0.075 39.825 897.870 55.397
room1 Mono Ours (Jetson) 0.803 0.00670 27.076 0.841 0.177 19.834 99.038 19.699
Ours (Laptop) 1.046 0.00868 30.459 0.902 0.092 21.580 333.430 32.959
Ours 1.183 0.00772 31.302 0.910 0.083 44.316 782.326 43.865
RGB-D Ours (Jetson) 0.381 0.00299 30.191 0.895 0.108 18.881 121.986 29.503
Ours (Laptop) 0.399 0.00277 33.071 0.931 0.062 21.782 367.455 43.568
Ours 0.394 0.00320 33.511 0.934 0.057 43.352 1018.111 49.617
room2 Mono Ours (Jetson) 0.241 0.00280 27.432 0.889 0.138 17.918 80.568 16.303
Ours (Laptop) 0.235 0.00263 32.970 0.935 0.075 19.499 339.442 22.790
Ours 0.225 0.00258 33.181 0.934 0.067 40.313 1053.078 26.834
RGB-D Ours (Jetson) 0.260 0.00257 31.883 0.928 0.078 15.982 95.480 33.592
Ours (Laptop) 0.275 0.00250 35.295 0.954 0.045 18.158 336.103 38.307
Ours 0.305 0.00257 35.028 0.951 0.043 36.244 953.755 41.032
Table 8: Detailed results of Photo-SLAM with different platforms on the Replica dataset.
8 More Results
The results of each scene of the replica dataset are detailed in Table 8 Additional qualitative results on the TUM dataset are demonstrated on Fig. 14 and Fig. 15, while Fig. 16 illustrates qualitative results of Photo-SLAM on EuReC Stereo Dataset.
Figure 14: Qualitative results of Photo-SLAM on TUM using monocular cameras.
Figure 15: Qualitative results of Photo-SLAM on TUM using RGB-D cameras.
Figure 16: Qualitative results of Photo-SLAM on stereo EuRoC.